OpenStudy (anonymous):

Find the arc length of the graph, on the interval [1/2,2] of y=(x^3/6)+(1/2x)

OpenStudy (ybarrap):

$$\large\mathbf \int^2_\frac{1}{2}\sqrt{1+ (\frac{dy}{dx}})^2 dx$$

OpenStudy (anonymous):

So which is that? 33/16 18/5 47/24 39/14